OpenStudy (anonymous):

Part A: The area of a square is (9x^2 - 12x + 4) square units. Determine the length of each side of the square by factoring the area expression completely. Show your work. (3 points) Part B: The area of a rectangle is (25x^2 - 16y^2) square units. Determine the dimensions of the rectangle by factoring the area expression completely. Show your work. (2 points) Part C: The volume of a rectangular box is (x^3 - 7x^2 - 9x + 63) cubic units. Determine the dimensions of the rectangular box by factoring the volume expression completely. Show your work. (5 points)

OpenStudy (campbell_st):

well the quadratic you have is a perfect square you need to factor it... to find the lengths the reason I say this is that \[9x^2, 4\] are both perfect squares, the middle term is double their product so the perfect square is \[a^2 -2ab + b^2 = (a - b)^2\] you just need to find a and b

OpenStudy (campbell_st):

B is the difference of 2 squares \[(a -b)(a +b) = a^2 = b^2\] find a and b and the last question can be do by starting with grouping in pairs \[(x^3 - 9x) - (7x^2 - 63)\] then factoring.. hope this all helps

OpenStudy (anonymous):

I agree with you, Campbell...I think you explained it perfectly and I dont think I have anything to add...great job! And good luck to you baby_sb

OpenStudy (anonymous):

I agree with you, Campbell...I think you explained it perfectly and I dont think I have anything to add...great job! And good luck to you baby_sb